OpenStudy (anonymous):

Find the slope and the y-intercept of the line C(1,1) and D(-2,3)

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

3-1/2-4 = -2/3

hartnn (hartnn):

x1-x2 = -2-1 = ?

OpenStudy (anonymous):

so,,, 1 = -2/3 (1) +b

OpenStudy (anonymous):

huh??? how did you get that?

hartnn (hartnn):

oh ,nothing, u mistyped it as 4, but your slope = -2/3 is correct :) and 1 = -2/3 (1) +b is also correct, find b from here

OpenStudy (anonymous):

b=3/2?

hartnn (hartnn):

b=1+2/3 = 5/3

OpenStudy (anonymous):

how so??

OpenStudy (anonymous):

the answer is 5/3,,, but i don't see how to get there from 1= -2/3(1) + b

hartnn (hartnn):

add 2/3 on both sides

OpenStudy (anonymous):

oh duh u add,,, i was multiplying

hartnn (hartnn):

1+2/3 = -2/3 +2/3 +b ..... yes.

OpenStudy (anonymous):

sigh,,,, ok tyvm!!!

hartnn (hartnn):

welcome ^_^